# Lennart Berg and Torben Wiedemann

### v1

@@ 8 The blueprint computation for $H_{3}$-graded groups
Unfortunately, the majority of the 15 blueprint identities are rather long and complicated: They fill a total of 20 A4 pages. Consequently, they are not accessible for manipulation by hand, so we have decided not to include them in this paper. However, since these identities hold for all possible choices of variables $y_{1},\ldots,y_{15}\in\mathcal{S}$, we can infer more succinct evaluated blueprint identities by replacing most of the indeterminates with $0_{\mathcal{S}}$ (or, in some cases, $1_{\mathcal{S}}$). The evaluated blueprint identities can be simplified by hand, which yields identities labelled (1), …, (35). We will not present how to deduce (1), …, (35) from the evaluated blueprint identities because this is a straightforward procedure, but we will give some remarks in the proofs of 8.7, 8.12 and 8.20. The interested reader can find the original blueprint identities, the evaluated blueprint identities and the code that produced them in [BW24].
@@
This follows from the second blueprint equation, which reads $-y_{14}-y_{5}y_{15}=-y_{14}-y_{15}y_{5}$. ∎
@@
for all $x,y\in\mathcal{S}$. This holds by 3.13.
@@
For all $\zeta,\xi\in H_{3}$ such that $\mathord{]\zeta,\xi[}$ contains exactly one element, the map $\psi_{\zeta,\xi}$ is bi-additive by 3.8.
@@
This is a straightforward computation, as explained in 8.2. For example, in order to derive (4), we put $y_{12}\coloneqq 1_{\mathcal{S}}$ and $y_{i}\coloneqq 0_{\mathcal{S}}$ for all $i\in\{1,\ldots,15\}\setminus\{4,5,12\}$ in blueprint identity 9. This yields the evaluated blueprint identity
@@
Using (2), (3) and 8.5, this identity is seen to be equivalent to (4). The identities have to be derived in the given order, so that each identity may be used in the proof of all subsequent identities. ∎
@@
Figure 7: Evaluation of the blueprint identities, part 1. See 8.7.
@@
Assertion i is (2). Assertion iii follows from (4), (5) and (6) and iv follows from (4) and (8). Assertion ii follows from (3) and (7) for the first three equalities and from (4) and 8.4 for the remaining equalities. By (10), (20) and (9), we have
@@
In particular, the images of the maps $\psi_{\alpha,\delta}^{\beta}$, $\psi_{\epsilon,\beta}^{\delta}$, $\psi_{\beta,\epsilon}^{\gamma}$, $\psi_{\delta,\alpha}^{\gamma}$ are identical, which allows us to apply (15) and (19) in a more general context. Assertion vi follows from 8.5, 8.6, (15), (19) and the previous observation. Assertion vii follows from (11), (13),  (8.11) and vi. Assertion v follows from (8.11), vi and vii. Finally, assertion viii is (21), assertion ix is (18) and assertion x is the culmination of (14) to (17). ∎
@@
We proceed as in 8.7. For most identities, the first step is to apply 8.10 x in order to remove zero summands. For (25), we use (24). For (30), we use (22) and (23). For (28), we use (24), (23) and (21). For (31), we use (21) and the substitutions $y_{5}\mapsto-y_{5}$ and $y_{15}\mapsto-y_{15}$. For (32), we use that the images of $f$ and $g$ are fixed by *. ∎
@@
Figure 8: Evaluation of the blueprint identities, part 2. See 8.12.
@@
Again, the map is additive by 8.10 vi. Further,
@@
By 8.13 and 8.15, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are ideals of $\mathcal{S}$. Further, for all $x\in\mathcal{S}$ and $y\in\mathcal{S}_{2}$, we have $g(x,y)=g(1,y)x=0x=0$ by (27).
@@
Figure 9: Evaluation of the blueprint identities, part 3. See 8.20.
@@
Let $x\in\mathcal{S}$. By 8.10 x, $f(x,1)$ lies in $\mathcal{S}_{2}$, so (33) yields that
@@
Hence $\operatorname{pr}_{1}(x)$ lies in $\mathcal{S}_{1}$. Further, $\operatorname{pr}_{2}(x)=-g(x,1)$ lies in $\mathcal{S}_{2}$ by 8.10 x.
@@
Since $g\big{(}1,h_{2}(x,1)\big{)}=-f\big{(}1,h_{2}(x,1)\big{)}=-g(x,1)=\operatorname{pr}_{2}(x)$ by 8.19 and 8.10 ix, we conclude that $x=\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)$. In particular, $\mathcal{S}=\mathcal{S}_{1}+\mathcal{S}_{2}$.
@@
by 8.19. Thus $x=\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)=0$. This finishes the proof. ∎
@@
It follows from 8.21 that $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ are nonassociative rings with identity elements $\operatorname{pr}_{1}(1_{\mathcal{S}})=-h_{1}(-f(1,1),1)$ and $\operatorname{pr}_{2}(1_{\mathcal{S}})=-g(1,1)=f(1,1)$, respectively, and that $\operatorname{pr}_{1}\colon\mathcal{S}\rightarrow\mathcal{S}_{1}$, $\operatorname{pr}_{2}\colon\mathcal{S}\rightarrow\mathcal{S}_{2}$ are homomorphisms.
Notation 8.23.
@@
The image of $\varphi$ lies in $\mathcal{S}_{2}$ by 8.10 x and it preserves the addition and multiplication by 8.13 and 8.22. Further, it follows from (24) that
@@
It is not true that the map $\mathcal{S}_{2}\rightarrow\mathcal{S}_{1},x\mapsto f(1,x)$ is an isomorphism. In fact, it is the zero map by 8.19.
@@
by 8.10 iv. Since $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$ is an isomorphism by 8.24, it follows that $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are associative. Hence $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is associative as well. ∎
@@
Let $x\in\mathcal{S}_{1}$. By 8.10 iii, we have $\varphi(x^{*})=f(1,x^{*})=-f(1,x)=\varphi(-x)$ and $\varphi(x)^{*}=f(1,x)^{*}=f(1,x)=\varphi(x)$. Since $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$ is an isomorphism and * is additive, the assertion follows. ∎
Notation 8.28.
@@
Put $x\coloneqq[x_{1},x_{2}],y\coloneqq[y_{1},y_{2}]\in\mathcal{S}$. By 8.10 x and 8.10 iv, we have
@@
Here $-f(y,1)x=-[0,y_{1}][x_{1},x_{2}]=[0,-y_{1}x_{2}]$. Since $h_{1}(\mathord{\>\cdot\>},1)=\varphi^{-1}$ on $\mathcal{S}_{2}$ by 8.24, we infer that
@@
Assertion iii follows. Finally, by 8.10 ix,
@@
The formulas in 8.29 show that the maps $f$, $g$, $h_{1}$, $h_{2}$ have the same form as the maps $\tilde{\psi}_{\alpha,\gamma}$, $\tilde{\psi}_{\alpha,\delta}^{\beta}$, $\tilde{\psi}_{\alpha,\varepsilon}^{\beta}$, $\tilde{\psi}_{\alpha,\varepsilon}^{\gamma}$ in Figure 4. Using 8.10, we infer that for all $\xi,\zeta\in\{\alpha,\beta,\gamma,\delta,\varepsilon\}$ and all $\rho\in\mathord{]\xi,\zeta[}$, the commutation map $\psi_{\xi,\zeta}^{\rho}$ in $G$ has the same form as the map $\tilde{\psi}_{\xi,\zeta}^{\rho}$ in Figure 4. A similar assertion holds for $\xi,\zeta\in\{\rho_{1},\rho_{1}+\rho_{2},\rho_{2}\}$ and for $\xi,\zeta\in\{\rho_{1},\rho_{3}\}$. Further, we have shown in  5.21 and 8.27 that the involution * has the same form as in the $H_{3}$-grading of $\operatorname{Ch}(D_{6},\mathcal{S}_{1})$ constructed in 4.10.
We can now state the main result of this paper, which will be reformulated in the language of foldings in 8.40. Note the distinction that we make between “parametrisations” and “coordinatisations”.
@@
Let $G$ be a group with an $H_{3}$-grading $(U_{\xi})_{\xi\in H_{3}}$ and fix a family $(w_{\rho})_{\rho\in\Delta}$ of Weyl elements in $G$. Denote by $\eta\colon H_{3}\times\Delta\rightarrow\{\pm 1\}^{2}$ the parity map from 6.16 and by $(\alpha,\beta,\gamma,\delta,\varepsilon)$ the $H_{2}$-quintuple corresponding to the $H_{2}$-pair $(\rho_{2},\rho_{3})=(\alpha,\varepsilon)$ in the fixed root base $\Delta$ of $H_{3}$. Then there exists a commutative ring $\mathcal{R}$ such that $G$ is coordinated by $\mathcal{R}\times\mathcal{R}$ with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$. By this we mean that there exists a parametrisation $(\theta_{\xi}\colon\mathcal{R}\times\mathcal{R}\rightarrow U_{\xi})_{\xi\in H_{3}}$ (in the sense of 6.9) of $(G,(U_{\xi})_{\xi\in H_{3}})$ by the standard parameter system ({±$(\{\pm 1\}^{2},\mathcal{R}\times\mathcal{R})$ for $\mathcal{R}$ from 6.10 with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$ such that the same commutator relations as in Figure 4 are satisfied.
@@
By 8.3 and 8.26, $\mathcal{R}\coloneqq\mathcal{S}_{1}$ is a commutative ring. By 8.1, 8.28 and 8.27, $(G,(U_{\alpha})_{\alpha\in H_{3}})$ is parametrised by the standard parameter system for $\mathcal{R}$ with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$. The desired commutator relations in the $H_{2}$-quintuple $(\alpha,\beta,\gamma,\delta,\varepsilon)$ hold by 8.30 and the commutator relation in the $A_{2}$-pair $(\rho_{1},\rho_{2})$ holds by the definition of the multiplication on $\mathcal{S}$ in 6.19. ∎
Remark 8.32.
Let everything be as in 8.31 and let $(\xi,\zeta)$ be a pair of non-proportional roots in $H_{3}$. Then there exists $w\in\operatorname{Weyl}(H_{3})$ such that $\xi^{w},\zeta^{w}$ lie in the positive part of the rank-2 subsystem spanned by two roots in $\Delta$. Thus we can conjugate one of the formulas in Figure 4 by an appropriate product of Weyl elements to obtain a commutator formula for $[\theta_{\xi}(r,s),\theta_{\zeta}(r^{\prime},s^{\prime})]$ where $r,s,r^{\prime},s^{\prime}\in\mathcal{R}$. We will see an example in 8.33.
It is a corollary of 8.31 that the $H_{3}$-grading $(U_{\alpha})_{\alpha\in H_{3}}$ of $G$ (and hence any $H_{3}$-grading of any group) is the folding of a $D_{6}$-grading in the sense of 4.10 (and 4.11). The remaining part of this paper deals with the technical details of this deduction. The first intermediate goal is 8.36.

### v2

@@ 9 The blueprint computation for $H_{3}$-graded
Unfortunately, the majority of the 15 blueprint identities are rather long and complicated: They fill a total of 20 A4 pages. Consequently, they are not accessible for manipulation by hand, so we have decided not to include them in this paper. However, since these identities hold for all possible choices of variables $y_{1},\ldots,y_{15}\in\mathcal{S}$, we can infer more succinct evaluated blueprint identities by replacing most of the indeterminates with $0_{\mathcal{S}}$ (or, in some cases, $1_{\mathcal{S}}$). The evaluated blueprint identities can be simplified by hand, which yields identities labelled (1), …, (35). We will not present how to deduce (1), …, (35) from the evaluated blueprint identities because this is a straightforward procedure, but we will give some remarks in the proofs of 9.7, 9.12 and 9.20. The interested reader can find the original blueprint identities, the evaluated blueprint identities and the code that produced them in [BW25].
@@
This follows from the second blueprint identity, which reads $-y_{14}-y_{5}y_{15}=-y_{14}-y_{15}y_{5}$. ∎
@@
for all $x,y\in\mathcal{S}$. This holds by 3.15.
@@
For all $\zeta,\xi\in H_{3}$ such that $\mathord{]\zeta,\xi[}$ contains exactly one element, the map $\psi_{\zeta,\xi}$ is bi-additive by 3.9.
@@
This is a straightforward computation, as explained in 9.2. For example, in order to derive (4), we put $y_{12}\coloneqq 1_{\mathcal{S}}$ and $y_{i}\coloneqq 0_{\mathcal{S}}$ for all $i\in\{1,\ldots,15\}\setminus\{4,5,12\}$ in blueprint identity 9. This yields the evaluated blueprint identity
@@
Using (2), (3) and 9.5, this identity is seen to be equivalent to (4). The identities have to be derived in the given order, so that each identity may be used in the proof of all subsequent identities. ∎
@@
Figure 7: Evaluation of the blueprint identities, part 1. See 9.7.
@@
Assertion i is (2). Assertion iii follows from (4), (5) and (6) and iv follows from (4) and (8). Assertion ii follows from (3) and (7) for the first three equalities and from (4) and 9.4 for the remaining equalities. By (10), (20) and (9), we have
@@
In particular, the images of the maps $\psi_{\alpha,\delta}^{\beta}$, $\psi_{\epsilon,\beta}^{\delta}$, $\psi_{\beta,\epsilon}^{\gamma}$, $\psi_{\delta,\alpha}^{\gamma}$ are identical, which allows us to apply (15) and (19) in a more general context. Assertion vi follows from 9.5, 9.6, (15), (19) and the previous observation. Assertion vii follows from (11), (13),  (9.11) and vi. Assertion v follows from (9.11), vi and vii. Finally, assertion viii is (21), assertion ix is (18) and assertion x is the culmination of (14) to (17). ∎
@@
We proceed as in 9.7. For most identities, the first step is to apply 9.10 x in order to remove zero summands. For (25), we use (24). For (30), we use (22) and (23). For (28), we use (24), (23) and (21). For (31), we use (21) and the substitutions $y_{5}\mapsto-y_{5}$ and $y_{15}\mapsto-y_{15}$. For (32), we use that the images of $f$ and $g$ are fixed by *. ∎
@@
Figure 8: Evaluation of the blueprint identities, part 2. See 9.12.
@@
Again, the map is additive by 9.10 vi. Further,
@@
By 9.13 and 9.15, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are ideals of $\mathcal{S}$. Further, for all $x\in\mathcal{S}$ and $y\in\mathcal{S}_{2}$, we have $g(x,y)=g(1,y)x=0x=0$ by (27).
@@
Figure 9: Evaluation of the blueprint identities, part 3. See 9.20.
@@
Let $x\in\mathcal{S}$. By 9.10 x, $f(x,1)$ lies in $\mathcal{S}_{2}$, so (33) yields that
@@
Hence $\operatorname{pr}_{1}(x)$ lies in $\mathcal{S}_{1}$. Further, $\operatorname{pr}_{2}(x)=-g(x,1)$ lies in $\mathcal{S}_{2}$ by 9.10 x.
@@
Since $g\big{(}1,h_{2}(x,1)\big{)}=-f\big{(}1,h_{2}(x,1)\big{)}=-g(x,1)=\operatorname{pr}_{2}(x)$ by 9.19 and 9.10 ix, we conclude that $x=\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)$. In particular, $\mathcal{S}=\mathcal{S}_{1}+\mathcal{S}_{2}$.
@@
by 9.19. Thus $x=\operatorname{pr}_{1}(x)+\operatorname{pr}_{2}(x)=0$. This finishes the proof. ∎
@@
It follows from 9.21 that $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ are nonassociative rings with identity elements $\operatorname{pr}_{1}(1_{\mathcal{S}})=-h_{1}(-f(1,1),1)$ and $\operatorname{pr}_{2}(1_{\mathcal{S}})=-g(1,1)=f(1,1)$, respectively, and that $\operatorname{pr}_{1}\colon\mathcal{S}\rightarrow\mathcal{S}_{1}$, $\operatorname{pr}_{2}\colon\mathcal{S}\rightarrow\mathcal{S}_{2}$ are homomorphisms.
Notation 9.23.
@@
The image of $\varphi$ lies in $\mathcal{S}_{2}$ by 9.10 x and it preserves the addition and multiplication by 9.13 and 9.22. Further, it follows from (24) that
@@
It is not true that the map $\mathcal{S}_{2}\rightarrow\mathcal{S}_{1},x\mapsto f(1,x)$ is an isomorphism. In fact, it is the zero map by 9.19.
@@
by 9.10 iv. Since $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$ is an isomorphism by 9.24, it follows that $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ are associative. Hence $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is associative as well. ∎
@@
Let $x\in\mathcal{S}_{1}$. By 9.10 iii, we have $\varphi(x^{*})=f(1,x^{*})=-f(1,x)=\varphi(-x)$ and $\varphi(x)^{*}=f(1,x)^{*}=f(1,x)=\varphi(x)$. Since $\varphi\colon\mathcal{S}_{1}\rightarrow\mathcal{S}_{2}$ is an isomorphism and * is additive, the assertion follows. ∎
Notation 9.28.
@@
Put $x\coloneqq[x_{1},x_{2}],y\coloneqq[y_{1},y_{2}]\in\mathcal{S}$. By 9.10 x and 9.10 iv, we have
@@
Here $-f(y,1)x=-[0,y_{1}][x_{1},x_{2}]=[0,-y_{1}x_{2}]$. Since $h_{1}(\mathord{\>\cdot\>},1)=\varphi^{-1}$ on $\mathcal{S}_{2}$ by 9.24, we infer that
@@
Assertion iii follows. Finally, by 9.10 ix,
@@
In the following proof, we apply the strategy from 7.2.
Proposition 9.30.
Assume that $\ell=4$. Then for all $x_{1},x_{2},y_{1},y_{2}\in\mathcal{S}_{1}$,
$[\theta_{\rho_{0}}([x_{1},x_{2}]),\theta_{\rho_{1}}([y_{1},y_{2}])]=\theta_{\rho_{0}+\rho_{1}}([x_{1}y_{1},x_{2}y_{2}]).$
Proof.
From the definition of the multiplication on $\mathcal{S}$ in (6.26), we already know that
$[\theta_{\rho_{1}}([x_{1},x_{2}]),\theta_{\rho_{2}}([y_{1},y_{2}])]=\theta_{\rho_{1}+\rho_{2}}([x_{1}y_{1},x_{2}y_{2}]).$
We want to conjugate this equation by $w_{\rho_{0}}w_{\rho_{1}}w_{\rho_{2}}$ because
$(\rho_{1},\rho_{2})^{\sigma((\rho_{0},\rho_{1},\rho_{2}))}=(\rho_{0}+\rho_{1},\rho_{2})^{\sigma((\rho_{1},\rho_{2}))}=(\rho_{0},\rho_{1}+\rho_{2})^{\sigma(\rho_{2})}=(\rho_{0},\rho_{1}).$
We can read off from Figures 12 and 13 that
$\eta(\rho_{1},(\rho_{0},\rho_{1},\rho_{2}))$ $=\eta(\rho_{1},\rho_{0})\eta(\rho_{0}+\rho_{1},\rho_{1})\eta(\rho_{0},\rho_{2})$
$=(-1,-1)(-1,-1)(1,1)=(1,1),$
$\eta(\rho_{2},(\rho_{0},\rho_{1},\rho_{2}))$ $=\eta(\rho_{2},\rho_{0})\eta(\rho_{2},\rho_{1})\eta(\rho_{1}+\rho_{2},\rho_{2})$
$=(1,1)(-1,-1)(-1,-1)=(1,1),$
$\eta(\rho_{1}+\rho_{2},(\rho_{0},\rho_{1},\rho_{2}))$ $=\eta(\rho_{1}+\rho_{2},\rho_{0})\eta(\rho_{0}+\rho_{1}+\rho_{2},\rho_{1})\eta(\rho_{0}+\rho_{1}+\rho_{2},\rho_{2})$
$=(-1,1)(1,1)(-1,1)=(1,1).$
Hence by conjugating the commutator relation above by $w_{\rho_{0}}w_{\rho_{1}}w_{\rho_{2}}$, we obtain the commutator relation in the assertion. ∎
@@
The formulas in 9.29 and 9.30 and the definition of the ring multiplication in (6.26) show that for all $\zeta,\xi\in E_{2}^{+}(\Delta)$ and all $\rho\in\mathord{]\zeta,\xi[}$, the commutation map $\psi_{\zeta,\xi}^{\rho}$ of $G$ is the same as in the folding of $\operatorname{Ch}(X,\mathcal{S}_{1})$ in 4.29. Further, it follows from 5.22 and 9.27 that the involution * on (the root groups of) $G$ has the same form as in $\operatorname{Ch}(X,\mathcal{S}_{1})$.
Proposition 9.32.
For all $\delta\in\Delta$, we have $w_{\delta}=\theta_{-\delta}(-1_{\mathcal{S}})\theta_{\delta}(1_{\mathcal{S}})\theta_{-\delta}(-1_{\mathcal{S}})$.
Proof.
By the definition of $1_{\mathcal{S}}$ in 6.24, we have $w_{\rho_{2}}=\theta_{-\rho_{2}}(a)\theta_{\rho_{2}}(1_{\mathcal{S}})\theta_{-\rho_{2}}(b)$ for some $a,b\in\mathcal{S}$. Here $a=b$ by  5.3 (iii) and
$\theta_{-\rho_{2}}(a)=\theta_{\rho_{2}}(1_{\mathcal{S}})^{w_{\rho_{2}}}=\theta_{\rho_{2}}(\eta(\rho_{2},\rho_{2}).1_{\mathcal{S}})$
by 5.3 (v). Since $\eta(\rho_{2},\rho_{2})=(-1,-1)$, it follows that $a=-1_{\mathcal{S}}$.
Put $w_{i}\coloneqq w_{\rho_{i}}$ for all $i\in\{0,1,2,3\}$ (the case $i=0$ only being allowed if $\ell=4$). For the remaining simple roots, we use the braid relations from 3.18. For example, it follows from $w_{1}w_{2}w_{1}=w_{2}w_{1}w_{2}$ that $w_{1}=w_{2}^{w_{1}w_{2}}$. Since $\eta(\rho_{2},(\rho_{1},\rho_{2}))=(1,1)=\eta(-\rho_{2},(\rho_{1},\rho_{2}))$, we infer that $w_{1}$ has the desired form. Similarly, $w_{0}=w_{1}^{w_{0}w_{1}}$ and $w_{0}$ has the desired form (if $\ell=4$). For $i=3$, we have $w_{2}w_{3}w_{2}w_{3}w_{2}=w_{3}w_{2}w_{3}w_{2}w_{3}$ and hence $w_{3}=w_{2}^{w_{3}w_{2}w_{3}w_{2}}$. As $\eta(\rho_{2},(\rho_{3},\rho_{2},\rho_{3},\rho_{2}))=(1,1)=\eta(-\rho_{2},(\rho_{3},\rho_{2},\rho_{3},\rho_{2}))$, we infer that $w_{3}$ has the desired form as well. ∎
We can now state the main result of this paper, which will be reformulated in the language of foldings in Section 10.
@@
Let $\ell\in\{3,4\}$, let $G$ be a group with an $H_{\ell}$-grading $(U_{\xi})_{\xi\in H_{\ell}}$ and fix a family $(w_{\rho})_{\rho\in\Delta}$ of Weyl elements in $G$. Then there exist a commutative ring $\mathcal{R}$ and a standard coordinatisation of $G$ by $\mathcal{R}\times\mathcal{R}$ with respect to $(w_{\rho})_{\rho\in\Delta}$ (in the sense of 6.28).
@@
By 9.3 and 9.26, $\mathcal{R}\coloneqq\mathcal{S}_{1}$ is a commutative ring. By 9.1, 9.28 and 9.27, $(G,(U_{\alpha})_{\alpha\in H_{3}})$ is parametrised by the standard parameter system for $\mathcal{R}$ with respect to $\eta$ and $(w_{\rho})_{\rho\in\Delta}$. The desired commutator relations hold by 9.31. Further, the Weyl elements $(w_{\rho})_{\rho\in\Delta}$ have the desired form by 9.32. ∎
10 All $H_{\ell}$-gradings arise from foldings
Notation 10.1.
In this section, we fix $\ell\in\{3,4\}$, an $H_{\ell}$-grading $(U_{\alpha}^{H})_{\alpha\in H_{\ell}}$ of a group $G$ and a family $(w_{\rho})_{\rho\in\Delta_{H}}$ of Weyl elements in $G$. We also choose a root base $\Delta_{H}$ of $H_{\ell}$ in standard order, which we denote by $(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ if $\ell=4$ and by $(\rho_{1},\rho_{2},\rho_{3})$ if $\ell=3$. As in 4.13, we denote by $\tilde{\pi}\colon X\rightarrow GH_{\ell}$, $\pi\colon X\rightarrow H_{\ell}$ the folding maps from 4.8 and by $\Delta_{X}=(\delta_{i})_{i=1}^{2\ell}$ the ordered root base of $X$ given by Figure 2. By 9.33, there exist a commutative ring $\mathcal{R}$ and a standard coordinatisation $(\theta_{\zeta}^{H})_{\zeta\in H_{\ell}}$ of $G$ by $\mathcal{R}\times\mathcal{R}$ with respect to $(w_{\rho})_{\rho\in\Delta_{H}}$, which we also fix. We further put $\mathcal{R}_{1}\coloneqq\mathcal{R}\times\{0\},\mathcal{R}_{2}\coloneqq\{0\}\times\mathcal{R}\subseteq\mathcal{R}\times\mathcal{R}$ and we denote elements $(x,y)$ of $\mathcal{R}\times\mathcal{R}$ by $[x,y]$, in analogy to 9.28.
It is a corollary of our main result 9.33 that the $H_{\ell}$-grading $(U_{\alpha}^{H})_{\alpha\in H_{\ell}}$ of $G$ (and hence any $H_{\ell}$-grading of any group) is the folding of an $X$-grading of $G$ in the sense of 4.19. This section deals with the technical details of this deduction. The first intermediate goal is 10.5.